# Dongyue Lu1, Lingdong Kong1, Gim Hee Lee1, Camille Simon Chane2, Wei Tsang Ooi1

### v1

@@ 3.2 FlexFuser: An Adaptive Event-Frame Fusion Module
${}^{(i)}\mathbf{h}_{\text{shared}}^{a}=\left[{}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a},\ ^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\right]\in\mathbb{R}^{B\times\left(C_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}+C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\right)\times H_{i}\times W_{i}}~{}.$ (4)
@@
Inspired by previous works (Zhou et al., 2023) and (Zhong et al., 2024), our goal is to dynamically fuse these two modalities in a flexible manner. The proposed FlexFuser module computes adaptive soft weights that regulate the contribution of each branch (event and frame) based on the current input conditions. As shown in Fig. 3, these adaptive soft weights are computed using a gating function, which incorporates learned noise to introduce perturbations for improved adaptability. The process is:
@@
$\left[{}^{(i)}\bm{\alpha},\ ^{(i)}\bm{\beta}\right]=\text{Softmax}\left((^{(i)}\mathbf{h}_{\text{shared}}^{a}\cdot^{(i)}\mathbf{W})+^{(i)}\sigma\cdot\epsilon\right),$ (5)
@@
where ${}^{(i)}\mathbf{W}\in\mathbb{R}^{\left(C_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}+C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\right)\times 2}$ is a trainable weight matrix, ${}^{(i)}\bm{\alpha}$ and ${}^{(i)}\bm{\beta}$ are the adaptive soft weights for the event and frame branches, respectively. Here, ${}^{(i)}\sigma$ is a learned standard deviation that controls the magnitude of the noise perturbation, and $\epsilon\sim\mathcal{N}(0,1)$ represents a Gaussian noise term.
The fused feature map at each scale $i$ is then obtained by applying the adaptive soft weights to the event and frame features:
@@
${}^{(i)}\mathbf{h}_{\text{fuse}}^{a}=^{(i)}\bm{\alpha}\odot^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a}+^{(i)}\bm{\beta}\odot^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}},$ (6)
@@
where $\odot$ denotes element-wise multiplication. This fusion process dynamically balances the contribution of each modality based on the input data, allowing for more robust and adaptive feature representation across varying conditions.
@@
Then, at each scale $i$, the final feature map combining event data at different frequencies and the frame data is obtained by adding the fused features from different frequencies. Specifically, we combine the fused feature maps as ${}^{(i)}\mathbf{h}_{\text{fuse}}=^{(i)}\mathbf{h}_{\text{fuse}}^{a}+^{(i)}\mathbf{h}_{\text{fuse}}^{b}$. After obtaining the fused feature maps across all scales, the multi-scale features are concatenated and fed into the detection head to produce the predicted bounding box $\mathbf{\hat{y}}$.
Optimization & Regularization. In addition to the standard detection loss $\mathcal{L}_{\text{det}}(\mathbf{y},\mathbf{\hat{y}})$, we introduce a regularization term to ensure balanced utilization of both the event and frame branches. This term penalizes large variations in the soft weights, encouraging a more uniform contribution from both modalities and preventing overfitting to a single branch:
@@
$\mathcal{L}_{\text{fuser}}=\mathcal{L}_{\text{det}}(\mathbf{y},\mathbf{\hat{y}})+\lambda\left(\frac{\text{Var}\left(\bm{\alpha}\right)}{\left(\mathbb{E}\left[\bm{\alpha}\right]\right)^{2}}+\frac{\text{Var}\left(\bm{\beta}\right)}{\left(\mathbb{E}\left[\bm{\beta}\right]\right)^{2}}\right),$ (7)
@@
3.3 FAL: Frequency-Adaptive Learning Mechanism
@@
FlexFuser aggregates information from different frequencies using labeled low-frequency data. To tune the model adaptively to handle diverse frequencies by leveraging both labeled low-frequency data and unlabeled high-frequency data, we design a flexible frequency-adaptive learning (FAL) mechanism. FAL incorporates multi-frequency information into the training process through iterative self-training. This approach enhances the model’s ability to generalize across varying frequencies, making it more robust in different scenarios. The key steps of the FAL mechanism are as follows:
@@
Pre-Training with Low-Frequency Labels. Rather than training solely at the data collection frequency $a$, we enhance the model’s capability by training it at a higher frequency $b$. To efficiently leverage the available labels, we select only the final event from the high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, which corresponds to the labeled timestamp. This approach allows the model to capture valuable high-frequency temporal information while still utilizing low-frequency labels, improving its temporal understanding and robustness. The training objective is to minimize the detection loss over the labeled data as follows:
$\mathcal{L}_{\text{GT}}=\sum_{(\mathbf{F},\mathbf{E}^{b}_{b/a},\mathbf{y})\in\mathcal{D}}\mathcal{L}_{\text{det}}(\mathbf{y},\mathbf{\hat{y}}).$ (8)
@@
Labels at Higher Frequencies. For the unlabeled data in $\mathcal{D}$ captured at frequency $b$, the pre-trained model generates high-frequency labels $\mathbf{\hat{y}}$ by performing inference on the entire high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$. These generated labels $\mathbf{\hat{y}}$ serve as pseudo-labels for guiding further training at higher frequencies, improving the model’s ability to generalize across different temporal conditions.
@@
Enhanced Temporal Refinement. To refine the high-frequency labels, we introduce a multi-step temporal refinement approach. First, we adopt bidirectional event augmentation by processing both forward and reversed event streams to detect objects with varying movements and orientations, thereby boosting recall. After generating the bidirectional labels, we apply Non-Maximum Suppression (NMS) to remove overlapping bounding boxes, followed by a low confidence threshold $\mathbf{\tau}$ to retain potential objects and further improve recall. Next, leveraging a tracking-by-detection framework, we link detection boxes across frames using pairwise IoU matching with a threshold $\mathbf{\tau}^{\text{IoU}}$. Short-lived tracks, with lengths below $\mathbf{L}^{\text{track}}$, are pruned to ensure temporal consistency. This approach ensures that the refined high-frequency labels $\tilde{\mathbf{y}}$ are accurate, temporally consistent, and reliable, ultimately improving detection quality in high-frequency data even in the absence of ground truth labels.
@@
Self-Training Iteration. The model is iteratively trained using the refined high-frequency labels $\tilde{\mathbf{y}}$ on high-frequency data where no ground truth labels are available. The total loss function combines the base training loss and the pseudo-label loss as:
$\mathcal{L}_{\text{FAL}}=\mathcal{L}_{\text{GT}}+\beta\sum_{(\mathbf{F},\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a-1},\mathbf{\tilde{y}})\in\mathcal{D}}\mathcal{L}_{\text{det}}\left(\mathbf{\tilde{y}},\mathbf{\hat{y}}\right)~{},$ (9)
@@
where $\beta$ balances the contribution of the high-frequency label loss. The complete FlexEvent framework combines FlexFuser and FAL, allowing the model to dynamically fuse event and frame data while adapting to varying frequencies. As demonstrated in the following sections, this combination provides a robust detection framework capable of maintaining high accuracy in dynamic environments.
@@ 4 Experiments
Table 1: Comparative study of state-of-the-art event camera detectors on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024). Both event-only and event-frame fusion methods are compared. The best and 2nd best scores from each metric are highlighted in bold and underlined, respectively.
Modality Method Venue Reference mAP AP50 AP75 APS APM APL
E RVT CVPR’23 (Gehrig & Scaramuzza, 2023) $38.4\%$ $58.7\%$ $41.3\%$ $29.5\%$ $50.3\%$ $81.7\%$
SAST CVPR’24 (Peng et al., 2024) $38.1\%$ $60.1\%$ $40.0\%$ $29.8\%$ $48.9\%$ $79.7\%$
SSM CVPR’24 (Zubić et al., 2024) $38.0\%$ $55.2\%$ $40.6\%$ $28.8\%$ $52.2\%$ $77.8\%$
LEOD CVPR’24 (Wu et al., 2024) $41.1\%$ $65.2\%$ $43.6\%$ $35.1\%$ $47.3\%$ $73.3\%$
E + F DAGr-18 Nature’24 (Gehrig & Scaramuzza, 2024) $37.6\%$
DAGr-34 Nature’24 (Gehrig & Scaramuzza, 2024) $39.0\%$
DAGr-50 Nature’24 (Gehrig & Scaramuzza, 2024) $41.9\%$ $66.0\%$ $44.3\%$ $36.3\%$ $56.2\%$ $77.8\%$
FlexEvent Ours - $\mathbf{57.4\%}$ $\mathbf{78.2\%}$ $\mathbf{66.6\%}$ $\mathbf{51.7\%}$ $\mathbf{64.9\%}$ $\mathbf{83.7\%}$
@@ 4.1 Experimental Settings
Datasets. We conduct experiments based on three large-scale event camera datasets: 1DSEC-Det (Gehrig & Scaramuzza, 2024), 2DSEC-Detection (Tomy et al., 2022), and 3DSEC-MOD (Zhou et al., 2023). These datasets comprise $78,344$ frames across $60$ sequences, $52,727$ frames over $41$ sequences, and $13,314$ frames within $16$ sequences, respectively, making them suitable for evaluating event-based object detection methods. We prioritize DSEC-Det as the primary benchmark for reporting results as it is the largest, most recent, and most comprehensive event-frame perception dataset. For more details, please refer to the Appendix.
Baselines. To evaluate our method, we compare it against both event-only and event-frame fusion state-of-the-art methods. For event-only detectors, we include RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), LEOD (Wu et al., 2024), and SSM (Zubić et al., 2024), retraining them on the DSEC-Det dataset following their respective protocols. For event-frame fusion methods, we compare with DAGr (Gehrig & Scaramuzza, 2024) on DSEC-Det using results from the original paper and retrain our method on DSEC-Detection and DSEC-MOD using standard settings to compare with CAFR (Cao et al., 2024) and RENet (Zhou et al., 2023). For other methods on DSEC-Detection and DSEC-MOD, we reference results reported in the CAFR and RENet papers. For more details, please refer to the Appendix.
Implementation Details. We train our model using the YOLOX framework (Zheng et al., 2021), optimizing with the standard detection loss, which consists of IoU loss, classification loss, and regression loss, averaged across both the batch and sequence length for stable training. Additionally, we introduce an extra regularization term to balance the utilization among the different modalities. The model is trained for $100,000$ iterations with a batch size of $8$ and a sequence length of $11$, using a learning rate of $1\times 10^{-4}$. All experiments are conducted on two NVIDIA RTX A5000 GPUs with 24GB memory, with the entire training process completed in approximately one day. Due to space limits, more details are included in the Appendix.
Evaluation Metrics. We evaluate object detectors using the mean Average Precision (mAP) as the primary metric, along with AP50, AP75, APS, APM, and APL from the COCO evaluation protocol (Lin et al., 2014). These metrics provide a comprehensive assessment of detection performance across different IoU thresholds and object sizes. Please refer to the Appendix for more details.
4.2 Comparisons to State-of-the-Art Detectors
Comparison to Event-Only Models. We compare FlexEvent with state-of-the-art event-only detectors, including RVT (Gehrig & Scaramuzza, 2023), SSM (Zubić et al., 2024), SAST (Peng et al., 2024), and LEOD (Wu et al., 2024), as shown in Tab. 1. We significantly outperform these methods across all metrics, with the performance gap becoming even more pronounced at higher frequencies. Event-only methods struggle to maintain detection accuracy at higher frequencies due to their inability to fully capture object semantics. In contrast, we overcome these limitations through the FlexFuser module, which integrates RGB data to compensate for the lack of semantic richness in the event stream. By fusing both event and frame data, we excel in complex and dynamic environments, achieving superior detection accuracy where event-only methods fall short.
@@ 4.2 Comparisons to State-of-the-Art Detectors
Table 2: Comparative study of state-of-the-art event camera detectors on the test set of DSEC-Detection (Tomy et al., 2022). Both event-only and event-frame fusion methods are compared. The reported results are the mAP scores of 1Car, 2Pedestrian (Ped), and 3Large-Vehicle (L-Veh) classes. The best and 2nd best scores from each metric are highlighted in bold and underlined, respectively.
Modality Method Venue Reference Type Car Ped L-Veh Average
E CAFR ECCV’24 (Cao et al., 2024) Event - - - $12.0\%$
E + F SENet CVPR’18 (Hu et al., 2018) Attention $38.4\%$ $14.9\%$ $26.0\%$ $26.2\%$
CBAM ECCV’18 (Woo et al., 2018) $37.7\%$ $13.5\%$ $27.0\%$ $26.1\%$
ECA-Net CVPR’20 (Wang et al., 2020) $36.7\%$ $12.8\%$ $27.5\%$ $25.7\%$
SAGate ECCV’20 (Chen et al., 2020) RGB + Depth $32.5\%$ $10.4\%$ $16.0\%$ $19.6\%$
DCF CVPR’21 (Ji et al., 2021) $36.3\%$ $12.7\%$ $28.0\%$ $25.7\%$
SPNet ICCV’21 (Zhou et al., 2021) $39.2\%$ $17.8\%$ $26.2\%$ $27.7\%$
RAMNet RA-L’21 (Gehrig et al., 2021a) $24.4\%$ $10.8\%$ $17.6\%$ $17.6\%$
FAGC Sensors’21 (Cao et al., 2021) $39.8\%$ $14.4\%$ $33.6\%$ $29.3\%$
FPN-Fusion ICRA’22 (Tomy et al., 2022) RGB + Event $37.5\%$ $10.9\%$ $24.9\%$ $24.4\%$
EFNet ECCV’22 (Sun et al., 2022a) $41.1\%$ $15.8\%$ $32.6\%$ $30.0\%$
DRFuser EAAI’23 (Munir et al., 2023) $38.6\%$ $15.1\%$ $30.6\%$ $28.1\%$
CMX TITS’23 (Zhang et al., 2023) $41.6\%$ $16.4\%$ $29.4\%$ $29.1\%$
RENet ICRA’23 (Zhou et al., 2023) $40.5\%$ $17.2\%$ $30.6\%$ $29.4\%$
CAFR ECCV’24 (Cao et al., 2024) $49.9\%$ $25.8\%$ $38.2\%$ $38.0\%$
FlexEvent Ours - $\mathbf{59.3\%}$ $\mathbf{37.4\%}$ $\mathbf{45.5\%}$ $\mathbf{47.4\%}$
Figure 4: Comparisons of event-based detectors under different event frequencies on DSEC-Det.
@@
Table 3: Comparisons of fusion-based event detectors on the test set of DSEC-MOD (Zhou et al., 2023). The best and 2nd best scores from each metric are highlighted in bold and underlined.
Method Venue Reference mAP
SENet CVPR’18 (Hu et al., 2018) $29.28\%$
CBAM ECCV’18 (Woo et al., 2018) $36.22\%$
ECA-Net CVPR’20 (Wang et al., 2020) $34.49\%$
SAGate ECCV’20 (Chen et al., 2020) $33.62\%$
DCF CVPR’21 (Ji et al., 2021) $32.20\%$
SPNet ICCV’21 (Zhou et al., 2021) $32.70\%$
FPN-Fusion ICRA’22 (Tomy et al., 2022) $32.28\%$
EFNet ECCV’22 (Sun et al., 2022a) $35.33\%$
RENet ICRA’23 (Zhou et al., 2023) $38.38\%$
FlexEvent Ours - $\mathbf{48.64\%}$
@@
Figure 5: Qualitative results of state-of-the-art event camera detectors. We compare FlexEvent with RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), and DAGr (Gehrig & Scaramuzza, 2024) on the test set of DSEC-Det. Best viewed in colors. See the Appendix for more examples.

### v3

@@ 3.2 FlexFuse: Adaptive Event-Frame Fusion
$^{(i)}\mathbf{h}_{\mathrm{shared}}^{a}=\left[{}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{a},\ ^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}\right]\in\mathbb{R}^{B\times\left(C_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}+C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}\right)\times H_{i}\times W_{i}}~.$ (4)
@@
Inspired by previous work [66, 64], our goal is to dynamically fuse these two modalities in a flexible manner. The proposed FlexFuse module computes adaptive soft weights that regulate the contribution of each branch (event and frame) based on the current input conditions. As shown in Fig. 3, these adaptive soft weights are computed using a gating function, which incorporates learned noise to introduce perturbations for improved adaptability:
@@
$^{(i)}\boldsymbol{\alpha},\ ^{(i)}\boldsymbol{\beta}]=\mathrm{Softmax}((^{(i)}\mathbf{h}_{\mathrm{shared}}^{a}\cdot^{(i)}\mathbf{W})+^{(i)}\sigma\cdot\epsilon),$ (5)
@@
where $^{(i)}\mathbf{W}\in\mathbb{R}^{\left(C_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}+C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}\right)\times 2}$ is a trainable weight matrix, $^{(i)}\boldsymbol{\alpha}$ and $^{(i)}\boldsymbol{\beta}$ are the adaptive soft weights for the event and frame branches, respectively. Here, $^{(i)}\sigma$ is a learned standard deviation that controls the magnitude of the noise perturbation, and $\epsilon\sim\mathcal{N}(0,1)$ represents a Gaussian noise term. The fused feature map at each scale $i$ is then obtained by applying adaptive soft weights:
@@
$^{(i)}\mathbf{h}_{\mathrm{fuse}}^{a}=^{(i)}\boldsymbol{\alpha}\odot^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{a}+^{(i)}\boldsymbol{\beta}\odot^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}},$ (6)
@@
where $\odot$ denotes element-wise multiplication. This fusion process dynamically balances the contribution of each modality based on the input data, allowing for more adaptive feature representation across varying conditions.
@@
Then, at each scale $i$, the final feature map combining event data at different frequencies and the frame data is obtained by adding the fused features from different frequencies. Specifically, we combine the fused feature maps as $^{(i)}\mathbf{h}_{\mathrm{fuse}}=^{(i)}\mathbf{h}_{\mathrm{fuse}}^{a}+^{(i)}\mathbf{h}_{\mathrm{fuse}}^{b}$. After obtaining the fused feature maps across all scales, the multi-scale features are concatenated and fed into the detection head to produce the predicted bounding box $\mathbf{\hat{y}}$.
[…]
@@
$\mathcal{L}_{\mathrm{fuse}}=\mathcal{L}_{\mathrm{det}}(\mathbf{y},\mathbf{\hat{y}})+\lambda\left(\frac{\mathrm{Var}\left(\boldsymbol{\alpha}\right)}{\left(\mathbb{E}\left[\boldsymbol{\alpha}\right]\right)^{2}}+\frac{\mathrm{Var}\left(\boldsymbol{\beta}\right)}{\left(\mathbb{E}\left[\boldsymbol{\beta}\right]\right)^{2}}\right),$ (7)
@@
3.3 FlexTune: Frequency-Adaptive Fine-Tuning
@@
FlexFuse aggregates information from different frequencies using labeled low-frequency data. To tune the model adaptively to handle diverse frequencies by leveraging both labeled low-frequency data and unlabeled high-frequency data, we design a FlexTune learning mechanism, incorporating multi-frequency information into the training process through iterative self-training. This enhances the ability to generalize across varying frequencies, making it more robust in different scenarios.
As shown in Fig. 4, FlexTune consists of two main stages: Low-Frequency Sparse Training to learn foundational knowledge from labeled data, and Cross-Frequency Propagation to transfer and refine the learned knowledge with high-frequency unlabeled data. This iterative mechanism bridges sparse supervision with dense temporal patterns, enabling robust detection under varying frequencies.
Figure 4: Illustration of the FlexTune learning mechanism. We first train on high-frequency events with sparse low-frequency labels, then we generate and refine high-frequency labels for cyclic self-training across frequencies.
@@
Low-Frequency Sparse Training. Rather than training solely at frequency $a$, we enhance the model’s capability by training it at a higher frequency $b$. To efficiently leverage the available sparse labels, we select only the last event from the high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, which corresponds to the labeled timestamp. This allows the model to capture valuable high-frequency temporal information while still utilizing low-frequency labels, improving its temporal understanding and robustness. The training objective is to minimize the detection loss over the sparse labeled data as:
[…]
where the summation is taken over samples $(\mathbf{F},\mathbf{E}^{b}_{b/a},\mathbf{y})\in\mathcal{D}$.
Cross-Frequency Propagation. Building on the low-frequency training, we transfer and refine model knowledge to unlabeled high-frequency data through three sequential steps: High-Frequency Bootstrapping, Temporal Consistency Calibration, and Cyclic Self-Training. Together, these steps generate accurate high-frequency pseudo-labels, enabling robust detection in diverse temporal settings.
@@
High-Frequency Bootstrapping. For the unlabeled data in $\mathcal{D}$ captured at frequency $b$, the pre-trained model generates high-frequency labels $\mathbf{\tilde{y}}$ by performing inference on the entire high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$. These generated labels $\mathbf{\tilde{y}}$ serve as pseudo-labels for bootstrapping further training at higher frequencies, improving the model’s ability to generalize across different temporal conditions.
@@
Temporal Consistency Calibration. To refine high-frequency labels, we enforce temporal coherence through three steps. (1) Bidirectional Event Augmentation. Process event streams forward and backward to capture diverse object motions, enhancing recall. (2) Confidence-Aware Filtering. Apply Non-Maximum Suppression (NMS) and a low confidence threshold $\mathbf{\tau}$ to eliminate duplicates and retain high-potential detections. (3) Tracklet Pruning. Link detections across frames using IoU-based tracking $(\mathbf{\tau}^{\mathrm{IoU}})$ and prune short tracks $(<\mathbf{L}^{\mathrm{track}})$ to suppress transient noise. This approach ensures that the refined high-frequency labels $\tilde{\mathbf{y}}$ are accurate, temporally consistent, and reliable, improving detection quality in high-frequency data even in the absence of ground truth labels.
@@
Cyclic Self-Training. The model is iteratively trained using the refined high-frequency labels $\tilde{\mathbf{y}}$ and the ground truth low-frequency label. The total loss function combines the base training loss and the pseudo-label loss as:
[…]
@@
where the summation here is taken over high-frequency samples $(\mathbf{F},\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a-1},\mathbf{\tilde{y}})\in\mathcal{D}$, and the coefficient $\beta$ balances the contribution of the high-frequency label loss. The complete FlexEvent framework combines FlexFuse and FlexTune, allowing the model to dynamically fuse event and frame data while adapting to varying frequencies.
@@ 4.1 Experimental Settings
Datasets. We conduct experiments based on three large-scale datasets: 1DSEC-Det [16], 2DSEC-Detection [50], and 3DSEC-MOD [66]. These datasets comprise 78,34478{,}344 frames across $60$ sequences, 52,72752{,}727 frames over $41$ sequences, and 13,31413{,}314 frames within $16$ sequences, respectively. making them suitable for evaluating event-based object detection methods. We prioritize DSEC-Det [16] as the primary benchmark for comparisons, as it is the largest, most recent, and most comprehensive event-frame perception dataset. We report results on Car and Pedestrian classes to ensure fair comparison with the state-of-the-art method DAGr [16]. For more details, please refer to Appendix Sec. A.1.
@@
Table 1: Comparative study of state-of-the-art event camera detectors on the validation set of the DSEC-Det [16] dataset. Both event-only and event-frame fusion methods are compared. The best and 2nd best scores from each metric are highlighted in bold and underlined, respectively.
   Modality Method Ref Venue mAP AP50 AP75 APS APM APL
E RVT [18] CVPR’23 $38.4$ $58.7$ $41.3$ $29.5$ $50.3$ $81.7$
SAST [39] CVPR’24 $38.1$ $60.1$ $40.0$ $29.8$ $48.9$ $79.7$
SSM [70] CVPR’24 $38.0$ $55.2$ $40.6$ $28.8$ $52.2$ $77.8$
LEOD [55] CVPR’24 $41.1$ $65.2$ $43.6$ $35.1$ $47.3$ $73.3$
E + F HDI-Former [31] arXiv’24 $46.7$ $69.1$ - - - -
DAGr-18 [16] Nature’24 $37.6$ - - - - -
DAGr-34 [16] Nature’24 $39.0$ - - - - -
     DAGr-50 [16]    Nature’24 $41.9$ $66.0$ $44.3$ $36.3$ $56.2$ $77.8$
FlexEvent - Ours $\mathbf{57.4}$ $\mathbf{78.2}$ $\mathbf{66.6}$ $\mathbf{51.7}$ $\mathbf{64.9}$ $\mathbf{83.7}$
@@
Table 2: Comparative study of state-of-the-art event camera detectors on the test set of DSEC-Detection [50]. Both event-only and event-frame fusion methods are compared. The reported results are the mAP scores of the 1Car, 2Pedestrian, and 3Large-Vehicle (L-Veh.) classes. The best and 2nd best scores from each metric are highlighted in bold and underlined, respectively.
   Modality Method Venue Ref Type   Car Pedestrian  L-Veh. Average
E CAFR    ECCV’24 [3] Event - - - $12.0$
SENet CVPR’18 [21] Attention  $38.4$ $14.9$ $26.0$ $26.2$
CBAM ECCV’18 [54]  $37.7$ $13.5$ $27.0$ $26.1$
ECA-Net CVPR’20 [53]  $36.7$ $12.8$ $27.5$ $25.7$
SAGate ECCV’20 [7] RGB + Depth  $32.5$ $10.4$ $16.0$ $19.6$
DCF CVPR’21 [24]  $36.3$ $12.7$ $28.0$ $25.7$
SPNet ICCV’21 [65]  $39.2$ $17.8$ $26.2$ $27.7$
RAMNet RA-L’21 [14] $24.4$ $10.8$ $17.6$ $17.6$
FAGC Sensors’21 [2] $39.8$ $14.4$ $33.6$ $29.3$
E + F    FPN-Fusion ICRA’22 [50]    RGB + Event  $37.5$ $10.9$ $24.9$ $24.4$
EFNet ECCV’22 [48] $41.1$ $15.8$ $32.6$ $30.0$
DRFuser EAAI’23 [37] $38.6$ $15.1$ $30.6$ $28.1$
CMX TITS’23 [58]  $41.6$ $16.4$ $29.4$ $29.1$
RENet ICRA’23 [66]  $40.5$ $17.2$ $30.6$ $29.4$
CAFR ECCV’24 [3]  $49.9$ $25.8$ $38.2$ $38.0$
FlexEvent Ours -  $\mathbf{59.3}$ $\mathbf{37.4}$ $\mathbf{45.5}$ $\mathbf{47.4}$
@@
Figure 5: Qualitative comparisons of state-of-the-art event camera detectors. We compare FlexEvent with RVT [18], SAST [39], and DAGr [16] on the test set of DSEC-Det. Best viewed in colors.
Figure 6: Results of detectors under varied frequencies on DSEC-Det.
Table 3: Comparisons of state-of-the-art event-frame fusion detectors on the test set of the DSEC-MOD [66] dataset.
Modal Method Venue Type F.mAP V.mAP Avg
E+F SENet [21] CVPR’18 Attention $29.3$ $17.5$ $23.4$
CBAM [54] ECCV’18 $36.2$ $16.4$ $26.3$
ECA-Net [53] CVPR’20 $34.5$ $18.8$ $26.7$
SAGate [7] ECCV’20 RGB + Depth $33.6$ $17.8$ $25.7$
DCF [24] CVPR’21 $32.2$ $18.8$ $25.5$
SPNet [65] ICCV’21 $32.7$ $14.7$ $23.7$
FPN-Fusion [50] ICRA’22 RGB + Event $32.3$ $15.0$ $23.7$
EFNet [48] ECCV’22 $35.3$ $18.1$ $26.7$
RENet [66] ICRA’23 $38.4$ $19.6$ $29.0$
FlexEvent Ours $\mathbf{48.6}$ $\mathbf{25.2}$ $\mathbf{36.9}$
Table 4: Comparative efficiency analysis of event detectors on DSEC-Det [16]. We report the inference times of event detectors at various frequencies, measured in milliseconds (ms).
Method Size Operational Frequency
(M) $\mathbf{20.0}$ Hz $\mathbf{36.0}$ Hz $\mathbf{90.0}$ Hz $\mathbf{180}$ Hz
RVT [18] $18.5$ $9.20$ ms $7.93$ ms $7.19$ ms $6.77$ ms
SAST [39] $18.9$ $14.06$ ms $12.37$ ms $11.52$ ms $11.10$ ms
SSM [70] $18.2$ $8.79$ ms $7.71$ ms $6.90$ ms $6.54$ ms
DAGr-50 [16] $34.6$ $73.35$ ms $55.11$ ms $45.29$ ms $43.89$ ms
FlexEvent $45.4$ $14.27$ ms $13.00$ ms $12.47$ ms $12.37$ ms